Toronto Blue Jays player George Springer to bond angle of 40 degrees from the horizontal towards the right field while on the trajectory of his ball will have it coming down at a point hundred 25 M from home plate and judge of the New York Yankees is position 30 metres from where the ball will come down and starts running at an instant the balls hit how fast would judge have to run to make the catch


1
Expert's answer
2021-03-08T08:23:25-0500
"R=\frac{v^2\sin{80}}{g}\\125=\frac{v^2\sin{80}}{9.8}\\v=35.3\frac{m}{s}"

"t=\frac{2v\sin{40}}{g}=\frac{2(35.3)\sin{40}}{9.8}=4.63\ s"

"u=\frac{30}{4.63}=6.5\frac{m}{s}"